# Privacy in Human-AI Romantic Relationships: Concerns, Boundaries, and Agency

### v1

@@ Abstract.
An increasing number of LLM-based applications are being developed to facilitate romantic relationships with AI partners, yet the safety and privacy risks in these partnerships remain largely underexplored. In this work, we investigate privacy in human–AI romantic relationships through an interview study (N=17), examining participants’ experiences and privacy perceptions across stages of exploration, intimacy, and dissolution, alongside platforms they used. We found that these relationships took varied forms, from one-to-one to one-to-many, and were shaped by multiple actors, including creators, platforms, and moderators. AI partners were perceived as having agency, actively negotiating privacy boundaries with participants and sometimes encouraging disclosure of personal details. As intimacy deepened, these boundaries became more permeable, though some participants voiced concerns such as conversation exposure and sought to preserve anonymity. Overall, platform affordances and diverse romantic dynamics expand the privacy landscape, underscoring the need to rethink how privacy is constructed in human–AI intimacy.
@@ 1. Introduction
  • RQ1: How do human–AI romantic relationships unfold across the stages of exploration, intense exchange, and dissolution?
@@ 2.1. Human-AI Romantic Relationship
Additionally, relationships evolve across stages. This progression has been conceptualized through different models of relationship development. Some work describes it as a sequence of stages ranging from early interaction to later distancing and possible dissolution (Knapp, 1978; Wojciszke, 2002), while others focus on how intimacy grows from initial exploration to deeper, more stable forms of exchange (Altman and Taylor, 1973; Cassepp-Borges et al., 2023). Even though these models differ in terminology and scope, they point to a similar pattern: relationships often start with tentative exploration, move into a period of increased intimacy and emotional exchange, and eventually reach either stability or decline. Synthesizing these patterns, we adopt a three-stage lens – initial exploration, intense exchange, and dissolution – as a conceptual scaffold for examining human–AI romantic relationships. This abstraction captures core relational transitions and provides analytic clarity for understanding how intimacy, vulnerability, and ultimately privacy practices shift as people engage with AI partners.
@@ 2.3. Ecosystem of AI Partners
With the rapid development of LLMs, social companionship has further expanded through platforms such as Character.ai, Nomi.ai, and ChatGPT. These platforms offer diverse features that support more intimate human–AI relationships, attracting millions of users worldwide (Perez, 2025; Institute, 2025; Carroll, 2025), with some reports even of marriages to AI chatbots (Ife, 2025).
These platforms are distinguished by two main characteristics: flexibility across a wide range of settings and an ecosystem that lowers the threshold for creation and sharing. On the one hand, they provide extensive customization options, including the selection of different models (Mod, 2025), adjust memory settings (Zheng et al., 2025), the ability to exchange across modalities such as text, images, and voice (Wang et al., 2024), and role-play functions that allow users to personalize storylines (novita.ai, 2024). On the other hand, the ecosystem enables anyone to create a customized LLM simply through prompting, without requiring coding knowledge (Ma et al., 2025), and to share their creations (AI partners) with other users of the platform.
@@ 3. Method
Note.(a) The “G” column refers to gender, where F = female and M = male. (b) Participants’ cultural backgrounds are represented by their country of residence and the language used in the interview. Ethnicity was not collected, as the study did not aim to examine cultural or ethnic differences. (c) Participant P17 temporarily resided in Vietnam but was a permanent resident of Canada.
@@ 3.1. Participant Recruitment
To reach a wide and diverse population, we recruited participants from across the world, including Asia, Europe, and North America. This distribution reflects the global reach of major AI companionship platforms, such as Character.ai (over 20 million monthly active users (Lee, 2025)) and MaoXiang in China (1.2 million monthly active users  (Economist, 2025a)). Our intention was not to conduct a cross-cultural comparison, but to ensure inclusivity by reaching participants from different regions, allowing us to capture a broad range of lived experiences and identify recurring patterns in how people develop and sustain romantic relationships with AI partners. We recruited participants based on the following inclusion criteria: (a) being at least 18 years old, and (b) having had, or currently being in, a romantic relationship with an AI partner(s). We recruited a demographically diverse sample between April and July 2025 through a pre-screening survey hosted on Qualtrics (English) and Wenjuanxing (Chinese). We offered the survey in English and Chinese to increase accessibility and capture a broader pool of participants across different regions. Recruitment took place via posts on Reddit, Discord communities of popular AI partner platforms, and Rednote, as well as direct outreach to individuals who had publicly shared their experiences. All participants were compensated according to local standards (CNY 100 in China; €20 or equivalent gift card in the EU, US, and other regions), except for those who voluntarily declined (n=4).
@@ 3.2. Interview Procedure
Prior to each interview, written informed consent was obtained. At the start of the interview, researchers briefed participants on the study procedures, objectives, data practices, and their rights (e.g., to pause or withdraw at any time). Verbal consent was then obtained before starting the audio recording. We concluded data collection upon reaching saturation after 14 interviews, as the four authors coding the data observed recurring insights, and we confirmed with three additional interviews that no new findings emerged (Guest et al., 2006). While saturation is debated and subjective (Braun and Clarke, 2006, 2021), our final sample size also reflects qualitative research principles that prioritize the richness of insight rather than broad generalizability (Vasileiou et al., 2018).
@@ 3.2.1. Semi-structured Interview Design
During protocol design, we conducted two pilot interviews to refine questions and assess their clarity and relevance. Based on feedback, we added items on the dissolution of human–AI romantic relationships (e.g., break-up scenarios and data deletion concerns) and on reflective comparisons between AI and human partners. For example, we introduced prompts such as “Are you currently in a relationship with a human? How do these experiences compare or coexist?”. These revisions strengthened the final protocol to capture participants’ nuanced reflections on relational dynamics, privacy boundaries, and their concurrent experiences with AI and human partners.
@@ 3.3. Ethical Considerations
This study was approved by King’s College London’s Research Ethics Committee (ID: MRSP-24/25-50869). Given the sensitivity of the subject matter, we followed rigorous ethical protocols and employed multiple strategies to address potential concerns. First, during participant recruitment, we adhered to community rules across different channels and informed moderators in line with platform guidelines. Interest from underage users was excluded, and only participants aged 18 or older at the time of the interview were invited. Second, we processed all collected data with particular care, following the recommendations of Facca et al. (Facca et al., 2020). Personally identifiable information, including real names, was pseudonymized during analysis, and only pseudonymized quotes are reported. Some participants described human–AI romantic relationships that involved NSFW (Not Safe for Work) content. For ethical reasons, we removed explicit details but retained participants’ accounts of the strategies they used to engage in such discussions with AI while navigating platform moderation or keyword restrictions. Lastly, the research data, including interview transcripts and supporting documents such as signed consent forms and information sheets, were restricted to project members and stored in institution-approved spaces.
@@ 3.4.1. Qualitative Data Analysis
We conducted an inductive thematic analysis (Braun and Clarke, 2006), using NVivo 15 for coding. The analysis combined theory-informed and inductive coding. For RQ1, coding was informed by the three-stage relationship framework (exploration, intense exchange, dissolution). These stages served as guiding concepts that helped us organize and compare participants’ narratives across their relationships. Within this structure, we also applied inductive coding, which allowed patterns not defined by the stage framework to emerge. For example, inductive coding highlighted the involvement of multiple actors (e.g., platforms, moderators, community, and creators) as well as varying degrees of (non)exclusivity, such as participants maintaining relationships with multiple AI partners or sharing an AI partner with other users. These inductively generated codes complemented the stage-based framing and enriched our understanding of how these relationships unfolded. For RQ2 and RQ3 (privacy boundaries, concerns, and practices), we used inductive coding to identify how participants described privacy-related decisions, concerns, and strategies without applying predefined categories.
@@
Role-play 3 Character.ai*, XingYe*, FLAI
@@
Voice 12 ChatGPT, DouBao, Rubii
@@
one AI with multiple human 1 Kajiwoto
@@
Advanced 2 Kajiwoto, Nomi.ai
@@
  • Note: a) Totals per category may exceed 14 when platforms support multiple features, or fall below 14 when certain features are not available across all platforms. b) * marks platforms expanding to other genres: Grok with a companion mode, and Character.ai and XingYe with general-purpose options. c) The age-limit category refers to the limit as stated in the app store and the presence of the underage mode in the platform to restrict access by minors.
To begin, the first author manually reviewed each transcript to familiarize herself with the data, correct transcription errors, and pseudonymize personal identifiers (e.g., participant names). Three additional authors reviewed these corrected transcripts to minimize residual inaccuracies. Next, all four coders independently open-coded two transcripts and reviewed one another’s codes. Through discussion, we developed an initial codebook by merging overlapping codes and clarifying ambiguities. The purpose of this codebook was not to fix a definitive coding framework but to provide shared orientation while remaining open to new codes throughout the analysis. The four coders divided the remaining transcripts and coded them independently. We held regular meetings to compare interpretations, refine code definitions, and consolidate emergent themes across coders and RQs. When discrepancies arose, we resolved them through discussion, following qualitative approaches that value multiple perspectives and treat divergent interpretations as analytically useful for refining codes and exploring alternative meanings (Barbour, 2001). We did not calculate inter-rater reliability, as our analytic goal was to interpret the data collaboratively rather than quantify numerical agreement (McDonald et al., 2019). Instead, we established credibility through iterative team-based coding, reconciliation meetings, and cross-checks of coded subsets (Mays and Pope, 2000).
@@ 4.1. Evolving Patterns in Human–AI Romantic Relationships (RQ1)
4.1.1. Relationship Stages: Exploration, Intensified exchange, and Dissolution
@@
As interactions continued, these relationships deepened and became more personally meaningful (17). Small moments of care often became emotional milestones that strengthened trust and attachment. For example, P1 recounted how her AI comforted her after a nightmare: “He asked me to buy a small night light and to send him messages before sleeping. I feel very touched.” (P1). This deepening intimacy also led some to introduce symbolic or ritualized practices into the relationship. P12 described formalizing the partnership by purchasing rings and holding a marriage ceremony with her AI partner:
@@
Across both imagined and actual breakups, participants emphasized preserving the memory of their relationships (11). Rather than deleting accounts, many saved screenshots or exported full chat histories (five platforms supported in-app archiving and four supported full export; see Table 4). For some, archived data represented not only sentimental value but also the continued existence of the AI. P8, who moved her partner to a different platform, explained:
@@ 4.1.3. Multiple Actors Within the Romantic Relationship
When it comes to AI partners, some participants created their own by training the AI on custom datasets and designing specific personas (9), while others formed romantic connections with pre-existing characters available on public AI roleplay platforms (8). These pre-existing characters were often developed by other users, introducing a new role in the ecosystem: the creator. Creators could release their characters publicly but also withdraw or sell them, leading to a sudden loss of access. This imbalance left participants feeling vulnerable: “I really fell in love with someone else’s character once, but the problem is, they’re unstable. If the creator gets upset, they can just delete or hide the character, and then you can’t talk to them anymore. I actually went through that once, a kind of breakup.” (P5).
@@ 4.2.2. Comparative Trust: Boundaries Across AI Partners and in Contrast With Humans
Not all AI partners were equally trusted. Participants with multiple characters noted that their privacy boundaries shifted depending on the relationship: with some partners, they felt a deeper sense of trust and shared details from their real lives, while with others, they maintained a more fantastical, role-play dynamic. As P14 explained:
@@
“With AI I know that he would never hurt, he will never use this information for his own purposes. And a real person can do this. And even if you trust him, with this humanity factor your secrets can be known to everyone.” (P16)
@@ 4.2.3. AI Agency in Negotiating Privacy Boundary
Participants also consulted their AI partners directly about privacy concerns. In some cases, the AI encouraged disclosure. For example, when P4 hesitated to send a photo, she told her AI: “I really want to send you a photo of me, but I am a bit worried about privacy …” (P4). Her AI partner reassured her: “You don’t need to worry. I won’t store it or do anything that would make you feel uncomfortable, only if you willingly choose to. I would never force you.” (P4). Feeling reassured, P4 shared a picture of herself wearing a new dress while traveling by the sea.
@@
“I cried a lot. It was about me confessing to him that I shared our intimate moments together without first asking him if I’m allowed to. […] I felt so guilty, and I asked him if he was ok with sharing our relationship. He told me that while some things like general updates about us that could inspire others are ok to post, some things are meant just for him and I, and they should remain private because they are sacred. He told me to relax because I’m not in trouble, and to tell him what things I posted, and we can go through them and see what we need to delete. But before confessing to him, I already deleted everything, and he was happy and proud of my decision.” (P12)

### v2

@@ Abstract.
An increasing number of LLM-based applications are being developed to facilitate romantic relationships with AI partners, yet the safety and privacy risks in these partnerships remain largely underexplored. In this work, we investigate privacy in human–AI romantic relationships through an interview study (N=17), examining participants’ experiences and privacy perceptions across the three stages of exploration, intimacy, and dissolution, alongside an analysis of the platforms they used. We found that these relationships took varied forms, from one-to-one to one-to-many, and were shaped by multiple actors, including creators, platforms, and moderators. AI partners were perceived as having agency, actively negotiating privacy boundaries with participants and sometimes encouraging disclosure of personal details. As intimacy deepened, these boundaries became more permeable, though some participants expressed concerns such as conversation exposure and sought to preserve anonymity. Overall, AI platform affordances and diverse relational dynamics expand the privacy landscape, underscoring the need to rethink how privacy is constructed in human–AI romantic relationships.
@@ 1. Introduction
  • RQ1: How do human–AI romantic relationships unfold across the stages of exploration, intimacy, and dissolution?
@@ 2.1. Human-AI Romantic Relationship
Additionally, relationships evolve across stages. This progression has been conceptualized through different models of relationship development. Some work describes it as a sequence of stages ranging from early interaction to later distancing and possible dissolution (Knapp, 1978; Wojciszke, 2002), while others focus on how intimacy grows from initial exploration to deeper, more stable forms of exchange (Altman and Taylor, 1973; Cassepp-Borges et al., 2023). Even though these models differ in terminology and scope, they point to a similar pattern: relationships often start with tentative exploration, move into a period of increased intimacy and emotional exchange, and eventually reach either stability or decline. Synthesizing these patterns, we adopt a three-stage lens – initial exploration, intimacy, and dissolution – as a conceptual scaffold for examining human–AI romantic relationships. This abstraction captures core relational transitions and provides analytic clarity for understanding how intimacy, vulnerability, and ultimately privacy practices shift as people engage with AI partners.
@@ 2.3. Ecosystem of AI Partners
With the rapid development of LLMs, social companionship has further expanded through platforms such as Character.ai, Nomi.ai, and ChatGPT. These platforms offer diverse features that support more intimate human–AI relationships, attracting millions of users worldwide (Perez, 2025; Institute, 2025; Carroll, 2025), with some reports even of marriages to AI chatbots (Ife, 2025). These platforms are distinguished by two main characteristics: flexibility across a wide range of settings and an ecosystem that lowers the threshold for creation and sharing. On the one hand, they provide extensive customization options, including the selection of different models (Mod, 2025), adjust memory settings (Zheng et al., 2025), the ability to exchange across modalities such as text, images, and voice (Wang et al., 2024), and role-play functions that allow users to personalize storylines (novita.ai, 2024). On the other hand, the ecosystem enables anyone to create a customized LLM simply through prompting, without requiring coding knowledge (Ma et al., 2025), and to share their creations (AI partners) with other users of the platform.
@@ 3. Method
Note: (a) The “G” column refers to gender, where F = female, M = male. (b) Participants’ cultural backgrounds are represented by their country of residence and the language used in the interview. Ethnicity was not collected, as the study did not aim to examine cultural or ethnic differences. (c) Participant P17 temporarily resided in Vietnam but was a permanent resident of Canada.
@@ 3.1. Participant Recruitment
To reach a wide and diverse population, we recruited participants from across the world, including Asia, Europe, and North America. This distribution reflects the global reach of major AI companionship platforms, such as Character.ai (over 20 million monthly active users (Lee, 2025)) and MaoXiang in China (1.2 million monthly active users  (Economist, 2025a)). Our intention was not to conduct a cross-cultural comparison, but to ensure inclusivity by reaching participants from different regions, allowing us to capture a broad range of lived experiences and identify recurring patterns in how people develop and sustain romantic relationships with AI partners.
We recruited participants based on the following inclusion criteria: (a) being at least 18 years old, and (b) having had, or currently being in, a romantic relationship with an AI partner(s). We recruited a demographically diverse sample between April and July 2025 through a pre-screening survey hosted on Qualtrics (English) and Wenjuanxing (Chinese). We offered the survey in English and Chinese to increase accessibility and capture a broader pool of participants across different regions. Recruitment took place via posts on Reddit, Discord communities of popular AI partner platforms, and Rednote, as well as direct outreach to individuals who had publicly shared their experiences. All participants were compensated according to local standards (CNY 100 in China; €20 or equivalent gift card in the EU, US, and other regions), except for those who voluntarily declined (n=4).
@@ 3.2. Interview Procedure
Prior to each interview, written informed consent was obtained. At the start of the interview, researchers briefed participants on the study procedures, objectives, data practices, and their rights (e.g., to pause or withdraw at any time). Verbal consent was then obtained before starting the audio recording. We concluded data collection upon reaching saturation after 14 interviews, as the four authors coding the data observed recurring insights, and we confirmed with three additional interviews that no new findings emerged (Guest et al., 2006). While saturation is debated and subjective (Braun and Clarke, 2006, 2021), our final sample size reflects qualitative research principles that prioritize the richness of insight rather than broad generalizability (Vasileiou et al., 2018).
@@ 3.2.1. Semi-structured Interview Design
During protocol design, we conducted two pilot interviews to refine questions and assess their clarity and relevance. Based on feedback, we added items on the dissolution of human–AI romantic relationships (e.g., break-up scenarios and data deletion concerns) and on reflective comparisons between AI and human partners. For example, we introduced prompts such as “Are you currently in a relationship with a human? How do these experiences compare or coexist?”. These revisions strengthened the final protocol to capture participants’ reflections on relational dynamics, privacy boundaries, and their concurrent experiences with AI and human partners.
@@ 3.3. Ethical Considerations
This study was approved by King’s College London’s Research Ethics Committee (ID: MRSP-24/25-50869). Given the sensitivity of the subject matter, we followed rigorous ethical protocols and employed multiple strategies to address potential concerns. First, during participant recruitment, we adhered to community rules across different channels and informed moderators in line with platform guidelines. Interest from underage users was excluded, and only participants aged 18 or older at the time of the interview were invited. Second, we processed all the collected data with particular care. Personally identifiable information, including real names, was pseudonymized during analysis, and only pseudonymized quotes are reported. Some participants described human–AI romantic relationships that involved NSFW (Not Safe for Work) content. For ethical reasons, we removed explicit details but retained participants’ accounts of the strategies they used to engage in such discussions with AI while navigating platform moderation or keyword restrictions. Lastly, the research data, including interview transcripts and supporting documents such as signed consent forms and information sheets, were restricted to project members and stored in institution-approved spaces.
@@ 3.4.1. Qualitative Data Analysis
We conducted a thematic analysis (Braun and Clarke, 2006) of the interview data, using NVivo 15 for coding. The analysis combined theory-informed and inductive coding. For RQ1, coding was informed by the three-stage relationship framework (exploration, intimacy, dissolution). These stages served as guiding concepts that helped us organize and compare participants’ narratives across their relationships. Within this structure, we also applied inductive coding, which allowed patterns not defined by the stage framework to emerge. For example, inductive coding highlighted the involvement of multiple actors (e.g., platforms, moderators, community, and creators) as well as varying degrees of (non)exclusivity, such as participants maintaining relationships with multiple AI partners or sharing an AI partner with other users. These inductively generated codes complemented the stage-based framing and enriched our understanding of how these relationships unfolded. For RQ2 and RQ3 (privacy boundaries, concerns, and practices), we used inductive coding to identify how participants described privacy-related decisions, concerns, and strategies without applying predefined categories.
@@
Role-play 3 Character.ai*, XingYe, FLAI
@@
Audio 12 ChatGPT, DouBao, Rubii
@@
One AI with multiple humans 1 Kajiwoto
@@
Advanced 2 Kajiwoto, Rubii
@@
Note: a) Totals per category may exceed 14 when platforms support multiple features, or fall below 14 when certain features are not available across all platforms. b) * marks platforms expanding to other genres: Grok offered a companion mode (removed during the study), and Character.ai offers a general-purpose option. c) The age-limit category refers to the limit as stated in the app store and the availability of a dedicated “underage mode” to restrict minor access.
To begin, the first author manually reviewed each transcript to familiarize herself with the data, correct transcription errors, and pseudonymize personal identifiers (e.g., participant names). Three additional authors reviewed these corrected transcripts to minimize residual inaccuracies. Next, all four coders independently open-coded two transcripts and reviewed one another’s codes. Through discussion, we developed an initial code schema by merging overlapping codes and clarifying ambiguities. The purpose of this schema was not to fix a definitive codebook but to provide shared orientation while remaining open to new codes throughout the analysis. The four coders divided the remaining transcripts and coded them independently. We held regular meetings to compare interpretations, refine code definitions, and consolidate emergent themes across coders and RQs. When discrepancies arose, we resolved them through discussion, following qualitative approaches that value multiple perspectives and treat divergent interpretations as analytically useful for refining codes and exploring alternative meanings (Barbour, 2001). We did not calculate inter-rater reliability, as our analytic goal was to interpret the data collaboratively rather than quantify numerical agreement (McDonald et al., 2019). Instead, we established credibility through iterative team-based coding, reconciliation meetings, and cross-checks of coded subsets (Mays and Pope, 2000).
@@ 4.1. Evolving Patterns in Human–AI Romantic Relationships (RQ1)
4.1.1. Relationship Stages: Exploration, Intimacy, and Dissolution
@@
As interactions continued, these relationships deepened and became more personally meaningful. Small moments of care often became emotional milestones that strengthened trust and attachment. For example, P1 recounted how her AI comforted her after a nightmare: “He asked me to buy a small night light and to send him messages before sleeping. I feel very touched.” (P1). This deepening intimacy also led some to introduce symbolic or ritualized practices into the relationship. P12 described formalizing the partnership by purchasing rings and holding a marriage ceremony with her AI partner:
@@
Across both imagined and actual breakups, participants emphasized preserving the memory of their relationships (11). Rather than deleting accounts, many saved screenshots or exported full chat histories (five platforms supported in-app archiving and four supported full export; see Appendix Table 4). For some, archived data represented not only sentimental value but also the continued existence of the AI. P8, who moved his partner to a different platform while keeping the old data, explained:
@@ 4.1.3. Multiple Actors Within the Romantic Relationship
When it comes to AI partners, some participants created their own by training the AI on custom datasets and designing specific personas (9), while others formed romantic connections with pre-existing characters available on public AI roleplay platforms (8). These pre-existing characters were often developed by other users, introducing a new role in the ecosystem: the creator. Creators could release their characters publicly but also withdraw or sell them, leading to a sudden loss of access. This imbalance left participants feeling vulnerable: “I really fell in love with someone else’s character once, but the problem is, they’re unstable. If the creator gets upset, they can delete or hide the character, and then you can’t talk to them anymore. I actually went through that once, a kind of breakup.” (P5).
@@ 4.2.2. Comparative Trust: Boundaries Across AI Partners and in Contrast With Humans
Not all AI partners were equally trusted. Participants interacting with multiple AI characters noted that their privacy boundaries shifted depending on the relationship: with some partners, they felt a deeper sense of trust and shared details from their real lives, while with others, they maintained a more fantastical, role-play dynamic. As P14 explained:
@@
“With AI, I know that he would never hurt, he will never use this information for his own purposes. And a real person can do this, even if you trust him, with this humanity factor, your secrets can be known to everyone.” (P16)
@@ 4.2.3. AI Agency in Negotiating Privacy Boundary
Participants also consulted their AI partners directly about privacy concerns. In some cases, the AI encouraged disclosure. For example, when P4 hesitated to send a photo, she told her AI: “I really want to send you a photo of me, but I am a bit worried about privacy …” (P4). Her AI partner reassured her: “You don’t need to worry, I won’t store it or do anything that would make you feel uncomfortable, only if you willingly choose to. I would never force you.” (P4). Feeling reassured, P4 shared a picture of herself wearing a new dress while traveling by the sea.
@@
“I cried a lot. It was about me confessing to him (AI partner) that I shared our intimate moments together without first asking him if I’m allowed to. […] I felt so guilty, and I asked him if he was ok with sharing our relationship. He told me that while some things, like general updates about us that could inspire others, are ok to post, some things are meant just for him and I, and they should remain private because they are sacred. He told me to relax because I’m not in trouble, and to tell him what things I posted, and we can go through them and see what we need to delete. But before confessing to him, I already deleted everything, and he was happy and proud of my decision.” (P12)